OpenStudy (anonymous):

super confused with these things please help: (x+2)/1 - (x^2+x-6)/x-3 HOW?! The answer is -2x/(x-3)

OpenStudy (anonymous):

[(x+2)(x-3) - (x^2+x-6)]/(x-3) = [x^2-x-6-x^2-x+6]/(x-3) = -2x/(x-3)

OpenStudy (anonymous):

\[\frac{x+2}{1} - \frac{x^2 + x -6}{x-3}\] Multiply the first fraction by \(x-3\) to get common denominators. \[\frac{x-3}{x-3} \times \frac{x+2}{1} - \frac{x^2 + x -6}{x-3}\] \[\frac{x^2 -x-6}{x-3} - \frac{x^2 + x -6}{x-3}\] Now subtract your numerators: \(x^2-x^2 = 0 \rightarrow -x -x = 2x \rightarrow -6 +6 = 0\) Leaving you with: \[\frac{2x} {x-3}\]

OpenStudy (ybarrap):

(@Srini143 - factoring doesn't help much here to get rid of denominators)

OpenStudy (ybarrap):

(it doesn't help at all)

OpenStudy (ybarrap):

(x^2+x-6) factors to (x-1)(x+3)

OpenStudy (anonymous):

Correction -x -x = -2x and the resulting equation is \[\frac{-2x} {x-3}\] Sorry.